# Methods for Legal Citation Prediction in the Age of LLMs: An Australian Law Case Study

### v1

@@ Abstract
In recent years, Large Language Models (LLMs) have shown great potential across a wide range of legal tasks. Despite these advances, mitigating hallucination remains a significant challenge, with state-of-the-art LLMs still frequently generating incorrect legal references. In this paper, we focus on the problem of legal citation prediction within the Australian law context, where correctly identifying and citing relevant legislations or precedents is critical. We compare several approaches: prompting general purpose and law-specialised LLMs, retrieval-only pipelines with both generic and domain-specific embeddings, task-specific instruction-tuning of LLMs, and hybrid strategies that combine LLMs with retrieval augmentation, query expansion, or voting ensembles. Our findings indicate that domain-specific pre-training alone is insufficient for achieving satisfactory citation accuracy even after law-specialised pre-training. In contrast, instruction tuning on our task-specific dataset dramatically boosts performance reaching the best results across all settings. We also highlight that database granularity along with the type of embeddings play a critical role in the performance of retrieval systems. Among retrieval-based approaches, hybrid methods consistently outperform retrieval-only setups, and among these, ensemble voting delivers the best result by combining the predictive quality of instruction-tuned LLMs with the retrieval system. [^1] For code, data, and models see https://auslawbench.github.io.
@@ 1 Introduction
Recent advancements in utilising Large Language Models (LLMs) for legal domain have shown promising results across various tasks. For instance, Pont et al. (2023) leveraged LLMs for generating summaries of judicial decisions, identifying legal issues, decision-making criteria, and specifying keywords. Deroy et al. (2024) revealed that LLMs outperform extractive summarisation methods in quality metrics but suffer from inconsistencies and hallucinations, highlighting the importance of human-in-the-loop approaches for improved reliability. Jiang & Yang (2023) underscored that fine-tuning LLMs demonstrates state-of-the-art performance in legal judgment prediction, while Peng & Chen (2024) showed retrieval-augmentation leads to improved accuracy by integrating external knowledge, particularly for complex charges. Hou et al. (2024) proposed a model to detect deviations between an AI-generated legal analysis and human as a way of quantifying their reliabilities. While these advancements highlight LLMs’ transformative potential in legal applications, challenges like ensuring factual accuracy, handling diverse tasks, and mitigating inherent issues such as hallucination remain. For instance, Dahl et al. (2024) reports that even state-of-the-art LLMs hallucinate between 69-88% of responses to legal queries, while Magesh et al. (2024) highlights that hallucination issue is mitigated with specialisation of tools but still persists as an unresolved issue.
In this paper, we report progress with respect to a specific task in the legal domain, Legal Citation Prediction, for Australian Law context. Citations in legal cases, as in academic writing, serve two purposes: first, information necessary to locate, read, and verify the material; and second, information about the authority of the source is conveyed (Axel-Lute, 1982). The second is particularly crucial in legal cases in common law jurisdictions, as in those systems most decision makers are required to follow previous decisions (Schauer, 1987). Accordingly, whilst precedent plays a fundamental role in determining how courts behave and, therefore, in how societies function, citations to, of, and between authorities are the way in which precedent is communicated. In determining matters in court, judges use citations to give weight and authority to their decisions and also, crucially, to demonstrate that they are acting appropriately and are adhering to the legal precedent that already exists. When doing so, a judge or a panel of judges will often state a legal rule – as a principal or a proposition – and then support the existence of that rule with a citation to another source; frequently a prior court decision. In other words, judges making determinations today rely on citations to not only determine legal questions in court but also to demonstrate that they are acting within their (the judges’) lawful role.
@@
In this paper, we aim to compare and explore different solutions for the citation prediction task. More specifically, we would like to make a comparison between the performance of the following approaches:
@@
  • Different hybrid tactics that combine LLMs and retrieval systems (i.e., retrieval-augmented generation, query expansion, voting ensemble)
2 Data and Methods
Data Source and Curation.
@@ 2 Data and Methods
We use the NSW Caselaw section of the Open Australian Legal Corpus  [^2] https://huggingface.co/datasets/umarbutler/open-australian-legal-corpus. We identified $82,530$ citations that specifically referred to a case within the dataset. For each citation ($c_{k}$), we extracted the sentence in which the citation appeared (denoted as $S^{i}_{c_{k}}$), along with the preceding sentence (denoted as $S^{i-1}_{c_{k}}$). We further utilised an LLM to generate an auxiliary description of $c_{k}$, based on (${{Full\ Text}_{c_{k}}},S^{i}_{c_{k}},S^{i-1}_{c_{k}}$) where ${Full\ Text}_{c_{k}}$ refers to the full text of the cited case. An example of a generated RoC is provided below:
@@
We refer to this description as Reason-of-Cite (RoC). [^3] In this process, we tasked the LLM to not generate any RoC (i.e., to generate NOT ENOUGH INFORMATION) if the information was not adequate in $(S^{i}_{c_{k}},S^{i-1}_{c_{k}})$. These refused cases were not included in the data. See Table A.3 of Appendix A.3 for details on when this was triggered. Each reference to a citation $c_{k}$ in the data results in a unique new $\text{RoC}_{c_{k}}$. For each $c_{k}$, we denote its $M$ references as $\mathrm{RoC}_{c_{k}}^{1},\mathrm{RoC}_{c_{k}}^{2},\ldots,\mathrm{RoC}_{c_{k}}^{M}$. We will discuss later how the RoCs are used. This resulted in the final dataset of $55,005$ instances, covering $18,677$ unique citations. Within this set, 5% of the citations have been referenced at least 9 times, while 54% were cited only once. From this final set, we extracted $1k$ citations as test set, and used the rest for training (i.e., our instruction-tuned models). Figure 1 (Top) shows the frequency distribution of all $18,677$ citations in the data, and Figure 1 (Bottom) shows the top-20 most frequent citations.
2.1 Methods
@@ 2.1.1 LLM-Only
For law-specialised LLMs, we used SaulLM-7B-instruct (Colombo et al., 2024b), SaulLM-54B-instruct (Colombo et al., 2024a) which to the best of our knowledge are the only publicly available law LLMs for English to this date. [^4] We were unable to run SaulLM-141B-Instruct due to its substantial hardware requirements. For prompting, we followed the exact setting of general purpose LLM experiments.
@@
We instruction-tuned LLaMA-3.1-8B and SaulLM-7B-Base models on the training data. These resulting models are referred to as Cite-LLaMA-3.1-8B and Cite-SaulLM-7B in our experiments. At inference time, given the query ($S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}})$) the model first produces the $\texttt{RoC}_{c_{k}}$ and then $c_{k}$ (i.e., $p(c_{k},\texttt{RoC}_{c_{k}}\mid I,S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}});\theta)$, where $\theta$ denotes LLM parameters, $I$ denotes the instruction, $\texttt{Mask}(S^{i}_{c_{k}})$ denotes $S^{i}_{c_{k}}$ with the citation to $c_{k}$ being masked). See Appendix A.2 for further details on training parameters and instruction detail and format. Unlike the previous setups, for all experiments with the instruction-tuned models only the text of the query (i.e., no RoC) was provided and the model is instructed to predict both the $RoC_{c_{k}}$ and the citation $c_{k}$. This presents a challenging setup as the LLM needs to predict a citation in the Open World setting solely by its parametric knowledge and the brief information provided in the query text.
2.1.2 Retrieval-Only
We investigated retrieval setup along two axes: embeddings, and granularity of data to be indexed in the vectorised database. The basic principle is a retrieval task where a query is matched against all records of a database, and the Top-k closest records are returned. The closeness is measured in the semantic space through vector embeddings. In all our setups involving retrieval, the search returns the Top-5 most relevant hits (i.e., citations).
@@ 2.1.2 Retrieval-Only
To construct the database, three possibilities for representing each citation ($c_{k}$) were considered: (1) Full Case Text ($\text{Full Case}_{c_{k}}$), (2) Catchwords ($\text{Catchwords}_{c_{k}}$), and (3) Aggregation of all its corresponding $M$ RoCs (RoC Aggregations=concat($\text{RoC Aggregations}=\operatorname{concat}(\mathrm{RoC}^{1}_{c_{k}},\mathrm{RoC}^{2}_{c_{k}},\ldots,\mathrm{RoC}^{M}_{c_{k}})$). We refer to these different settings as Index Granularity later in the experiments (i.e., Table 1). For each granularity level, we created a distinct database version for each embedding backbone.  [^8] We use https://github.com/chroma-core/chroma for building the DB, and use the built-in cosine similarity for measuring closeness.
2.1.3 Hybrid of LLM and Retrieval
@@
This follows a standard Retrieval-Augmented Generation (RAG) setup, where queries are sent to the database, and the Top-5 results are retrieved for the LLM to re-rank and select the Top-1 match.
@@ 3 Experiments
Closed World Embeddings Index Granularity Query Output ACC@1 ACC@5
@@
Closed World text-embedding-3-large Full Cases Text Top-5 Citations 14.3/14.4/17.1/17.4 31.1/31.4/34.3/34.4
@@
Closed World Embeddings Index Granularity Query Output ACC@1 ACC@5
@@
Table 1: Citation accuracy under various methods, including ours. For results marked by $\ddagger$, we used gpt-4o-mini due to cost. Best result is marked with . The rows are colour-coded for readability and easier comparison of settings referenced in the discussion §3.1. Text denotes ($S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}})$), see §2.1.1 for notations.
Experimental Setups.

### v2

# Evaluating LLM-based Approaches to Legal Citation Prediction: Domain-specific Pre-training, Fine-tuning, or RAG? A Benchmark and an Australian Law Case Study

@@ Abstract
Large Language Models (LLMs) have demonstrated strong potential across legal tasks, yet the problem of legal citation prediction remains under-explored. At its core, this task demands fine-grained contextual understanding and precise identification of relevant legislation or precedent. We introduce the AusLaw Citation Benchmark, a real-world dataset comprising $55k$ Australian legal instances and $18,677$ unique citations which to the best of our knowledge is the first of its scale and scope. We then conduct a systematic benchmarking across a range of solutions: (i) standard prompting of both general and law-specialised LLMs, (ii) retrieval-only pipelines with both generic and domain-specific embeddings, (iii) supervised fine-tuning, and (iv) several hybrid strategies that combine LLMs with retrieval augmentation through query expansion, voting ensembles, or re-ranking. Results show that neither general nor law-specific LLMs suffice as stand-alone solutions, with performance near zero. Instruction tuning (of even a generic open-source LLM) on task-specific dataset is among the best performing solutions. We highlight that database granularity along with the type of embeddings play a critical role in retrieval-based approaches, with hybrid methods which utilise a trained re-ranker delivering the best results. Despite this, a performance gap of nearly $50\%$ remains, underscoring the value of this challenging benchmark as a rigorous test-bed for future research in legal-domain. [^1] The data, code, and trained LLMs and re-rankers are available at: https://auslawbench.github.io
Evaluating LLM-based Approaches to Legal Citation Prediction: Domain-specific Pre-training, Fine-tuning, or RAG?
A Benchmark and an Australian Law Case Study
Jiuzhou Han     Paul Burgess     Ehsan Shareghi  Department of Data Science & AI, Monash University  Faculty of Law, Monash University first.last@monash.edu
@@ 1 Introduction
Recent advancements in utilising Large Language Models (LLMs) for legal domain have shown promising results across various tasks. For instance, Pont et al. (2023) leveraged LLMs for generating summaries of judicial decisions, identifying legal issues, decision-making criteria, and specifying keywords. Deroy et al. (2024) revealed that LLMs outperform extractive summarisation methods in quality metrics but suffer from inconsistencies and hallucinations, highlighting the importance of human-in-the-loop approaches for improved reliability. Jiang and Yang (2023) underscored that fine-tuning LLMs demonstrates state-of-the-art performance in legal judgment prediction, while Peng and Chen (2024) showed retrieval-augmentation leads to improved accuracy by integrating external knowledge, particularly for complex charges. Hou et al. (2024a) proposed a model to detect deviations between an AI-generated legal analysis and human as a way of quantifying their reliabilities.
While these advancements highlight LLMs’ transformative potential in legal applications, challenges like ensuring factual accuracy, handling diverse tasks, and mitigating inherent issues such as hallucination remain. For instance, Dahl et al. (2024) reports that even state-of-the-art LLMs hallucinate between 69-88% of responses to legal queries, while Magesh et al. (2024) highlights that hallucination issue is mitigated with specialisation of tools but still persists as an unresolved issue.
In this paper, we report progress with respect to a less-explored task in the legal domain, Legal Citation Prediction. Citations in legal cases, as in academic writing, serve two purposes: first, information necessary to locate, read, and verify the material; and second, information about the authority of the source is conveyed (Axel-Lute, 1982). The second is particularly crucial in legal cases in common law jurisdictions, as in those systems most decision makers are required to follow previous decisions (Schauer, 1987). Accordingly, whilst precedent plays a fundamental role in determining how courts behave and, therefore, in how societies function, citations to, of, and between authorities are the way in which precedent is communicated. In determining matters in court, judges use citations to give weight and authority to their decisions and also, crucially, to demonstrate that they are acting appropriately and are adhering to the legal precedent that already exists. When doing so, a judge or a panel of judges will often state a legal rule – as a principal or a proposition – and then support the existence of that rule with a citation to another source; frequently a prior court decision. In other words, judges making determinations today rely on citations to not only determine legal questions in court but also to show that they are acting within their (the judges’) lawful role.
@@
We aim to compare, develop, and explore different solutions for the citation prediction task. We release, AusLaw Citation Benchmark, a real dataset of $55k$ instances specific to Australian law, covering $18,677$ unique citations. To the best of our knowledge this is the first dataset of this scale and scope. We then conduct a thorough comparison along the following comprehensive dimensions:
@@
  • Different hybrid tactics that combine LLMs and retrieval systems (i.e., retrieval-augmented generation, query expansion, voting ensemble, and specialised re-rankers)
While we aim to push all these solutions to their limits to better understand their effectiveness and limitations, a near $50\%$ gap is remained to be filled. We hope this benchmark and the baselines and methods investigated in this paper to encourage future developments in this important and challenging task.
2 Legal NLP in the Era of LLMs
While a satisfying review demands a separate work, we attempt to provide a brief overview of existing work on the intersection of Law and LLMs. We also acknowledge that this only provides an overview of a small subset of the broader space of research in Legal NLP.
The intersection of NLP and the legal domain has given rise to a wide array of research tackling tasks such as legal text classification, retrieval, summarisation, question answering, and reasoning. As legal texts are often complex and domain-specific, specialised benchmarks and models have become essential to evaluate and improve NLP performance in this high-stakes field.
2.1 Legal Benchmarks
Recent efforts in benchmarking legal NLP have produced a diverse ecosystem of datasets targeting legal reasoning, retrieval, and comprehension. Benchmarks like LEGALBENCH Guha et al. (2024), LawBench Fei et al. (2023), and LexEval offer broad, multi-task evaluations across cognitive levels and languages, revealing persistent limitations in current LLMs’ legal reasoning abilities. Task-specific benchmarks like BLT Blair-Stanek et al. (2023) and MAUD Wang et al. (2023) assess fundamental skills like basic legal text navigation and merger agreement comprehension, while CUAD Hendrycks et al. (2021) and the statutory reasoning Holzenberger et al. (2020) in tax law dataset bring attention to contract review and rule-based reasoning tasks demanding deep understanding of legal semantics and structure. LexGLUE Chalkidis et al. (2022) and LEGAL-BERT Chalkidis et al. (2020) contribute to general legal language understanding and adaptation of pretrained models for legal tasks, while LegalBench-RAG Pipitone and Alami (2024) and CLERC Hou et al. (2024b) evaluate retrieval and generation in legal writing tasks. Newer benchmarks such as LegalAgentBench Li et al. (2024a) assess autonomous legal agents through multi-step tasks and tool use in real-world scenarios, while LegalHalBench Hu et al. (2025) introduces fine-grained metrics and datasets to evaluate and reduce hallucination in legal QA systems. Region-specific datasets like IL-TUR Joshi et al. (2024) for Indian law and LexEval Li et al. (2024b) for Chinese legal systems underscore the need for jurisdictionally relevant evaluations. We propose a new benchmark focused on legal citation prediction.
2.2 Law-Specialised Large Language Models
The development of law-specialised LLMs has gained momentum in response to the distinctive demands of legal language, which is domain-specific, syntactically complex, and semantically nuanced. Early adaptation efforts such as LEGAL-BERT Chalkidis et al. (2020) explored various strategies for pretraining BERT models on legal corpora, revealing that full domain-specific pretraining outperforms generic approaches. More recent models like Lawma Dominguez-Olmedo et al. (2024) and LawLLM Shu et al. (2024) demonstrate the effectiveness of fine-tuned and multi-task architectures in addressing legal classification, retrieval, and judgment prediction tasks within the United States legal system. The SaulLM Colombo et al. (2024b, a) family is the first (and only) open-source Law-specialised LLMs ranging from 7B to 141B parameters, using extensive legal corpora (500B tokens), instruction tuning, and preference alignment to achieve state-of-the-art results across multiple legal benchmarks. These models highlight the benefits of legal specialisation.
Nevertheless, the presence of domain-specialised LLMs only reduce the semantic gap required for the domain and still require (as we will show later in the experiments) task-specific training to be competent for fine-grained tasks such as legal citation prediction. Additionally, we argue (and empirically show) that pretraining on a mixture of data from various jurisdictions and countries does not provide a reliable understanding of the legal context specific to each of those jurisdictions. To bridge this gap, we develop an Australia-specific legal LLM focused on tasks such as legal citation prediction—thereby supporting jurisdiction-aware applications and advancing the broader field of legal artificial intelligence.
3 AusLaw Citation Benchmark
@@
To build our benchmark, we use the NSW Caselaw section of the Open Australian Legal Corpus. [^2] https://huggingface.co/datasets/umarbutler/open-australian-legal-corpus We identified $82,530$ citations that specifically referred to a case within the dataset. For each citation ($c_{k}$), we extracted the sentence in which the citation appeared (denoted as $S^{i}_{c_{k}}$), along with the preceding sentence (denoted as $S^{i-1}_{c_{k}}$). We further utilised an LLM to generate an auxiliary description of $c_{k}$, based on (${{Full\ Text}_{c_{k}}},S^{i}_{c_{k}},S^{i-1}_{c_{k}}$) where ${Full\ Text}_{c_{k}}$ refers to the full text of the cited case. We manually checked a subset of the LLM-generated descriptions for quality assurance and optimizing the prompt wording. An example of a generated RoC is provided below:
@@
We refer to this description as Reason-of-Citation (RoC). [^3] In this process, we tasked the LLM to not generate any RoC (i.e., to generate NOT ENOUGH INFORMATION) if the information was not adequate in $(S^{i}_{c_{k}},S^{i-1}_{c_{k}})$. These refused cases were not included in the data. See Table A.4 of Appendix A.4 for details on when this was triggered. Each reference to a citation $c_{k}$ in the data results in a unique new $\text{RoC}_{c_{k}}$. For each $c_{k}$, we denote its $M$ references as $\mathrm{RoC}_{c_{k}}^{1},\mathrm{RoC}_{c_{k}}^{2},\ldots,\mathrm{RoC}_{c_{k}}^{M}$. We will discuss later how the RoCs are used. This resulted in the final dataset of $55,005$ instances, covering $18,677$ unique citations. Within this set, 5% of the citations have been referenced at least 9 times, while 54% were cited only once. From this final set, we extracted $1k$ citations as test set, and used the rest for training (i.e., our instruction-tuned models). See Appendix A.2 for more details.
4 Methods
@@ 4.1 LLM-Only
For law-specialised LLMs, we used SaulLM-7B-instruct (Colombo et al., 2024b), SaulLM-54B-instruct (Colombo et al., 2024a) which to the best of our knowledge are the only publicly available law LLMs for English to this date. [^4] We were unable to run SaulLM-141B-Instruct due to its substantial hardware requirements. For prompting, we followed the exact setting of general purpose LLM experiments (see Table A.4 of Appendix A.4).
@@
We instruction-tuned LLaMA-3.1-8B and SaulLM-7B-Base models on the training data. These resulting models are referred to as Cite-LLaMA-3.1-8B and Cite-SaulLM-7B in our experiments. At inference time, given the query ($S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}})$) the model first produces the $\texttt{RoC}_{c_{k}}$ and then $c_{k}$ (i.e., $p(c_{k},\texttt{RoC}_{c_{k}}\mid I,S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}});\theta)$:
$p(c_{k}\mid I,S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}}),\texttt{RoC}_{c_{k}};\theta)$ (1)
×$\quad\times p(\texttt{RoC}_{c_{k}}\mid I,S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}});\theta).$
where $\theta$ denotes LLM parameters, $I$ denotes the instruction, $\texttt{Mask}(S^{i}_{c_{k}})$ denotes $S^{i}_{c_{k}}$ with the citation to $c_{k}$ being masked. See Appendix A.3 for further details on training parameters and instruction detail and format (same instruction was used at inference).
Unlike the previous setups, for all experiments with the instruction-tuned models only the text of the query (i.e., no RoC) was provided and the model is instructed to predict both the $RoC_{c_{k}}$ and the citation $c_{k}$. This presents a challenging setup as the LLM needs to predict a citation in the Open World setting solely by its parametric knowledge and the brief information provided in the query text.
4.2 Retrieval-Only
We investigated retrieval along two axes: embeddings, and granularity of data to be indexed in the vectorised database. The basic principle is a retrieval task where a query is matched against all records of a database, and the Top-k closest records are returned. The closeness is measured in the semantic space through embeddings. In all setups involving retrieval, unless stated otherwise, the search returns the Top-5 most relevant citations.
@@ 4.2 Retrieval-Only
To construct the database, three possibilities for representing each citation ($c_{k}$) were considered: (1) Full Case Text ($\text{Full Case}_{c_{k}}$), (2) Catchwords ($\text{Catchwords}_{c_{k}}$), and (3) Aggregation of all its corresponding $M$ RoCs (RoC Aggregations=concat($\text{RoC Aggregations}=\operatorname{concat}(\mathrm{RoC}^{1}_{c_{k}},\mathrm{RoC}^{2}_{c_{k}},\ldots,\mathrm{RoC}^{M}_{c_{k}})$).
We refer to these different settings as Index Granularity later in the experiments (i.e., Table 1). For each granularity level, we created a distinct database version for each embedding backbone. [^8] We use https://github.com/chroma-core/chroma for building the DB, and use the built-in cosine similarity for measuring closeness.
4.3 Hybrid of LLM and Retrieval
@@
This follows a standard Retrieval-Augmented Generation (RAG) setup, where queries are sent to the database, and the Top-5 results are retrieved for the LLM to re-rank and select the Top-1 match. See Table A.4 of Appendix A.4 for prompt details.
Re-ranker.
We have four settings to train the re-ranker. The same data used for instruction fine-tuning was used for training the re-rankers. The input of a re-ranker is the text containing the missing citation and retrieved top-5 candidate citations and their corresponding RoC. The output is the gold citation. Since the RoC of each case has different length, to train a re-ranker, we needed to eliminate the effect of discrepancy in the length.
  • Setting 1 uses the first reason as the whole citation reason.
  • Setting 2 tasks GPT-4o-mini to merge the citation reasons of each case into a similar length text.
  • Setting 3 leverages the same merged citation reason as Setting 2. In addition, given 5 cases and their case texts, the GPT-4o-mini is tasked to generate the rational of why the gold one is cited among all 5 citations.
  • Setting 4 is based on retrieval. We retrieve the most similar citation reason from the citation reasons list as the citation reason of that case. Within this setting, there are two variations: 1) we use the retrieved top-1 citation reason as the output for the training; 2) we use the gold citation reason as the output for the training. We name these two variations as 4v1 and 4v2.
See Table 10 of Appendix A.3 for input and output details.
@@ 5 Experiments
Embeddings Index Granularity Query Output ACC@1 ACC@5
@@
text-embedding-3-large Full Cases Text Top-5 Citations 14.3/14.4/17.1/17.4 31.1/31.4/34.3/34.4
@@
Embeddings Index Granularity Query Output ACC@1 ACC@5
@@
         (Hybrid Approach) Re-ranker: Given the Query, first retrieves Top-5 citations, and use trained re-ranker to pick the best
Results are formatted as setting1/2/3/4v1/4v2
Embeddings Index Granularity Query Output ACC@1 ACC@5
text-embedding-3-large RoC Aggregations Text Top-5 Citations 20.2/26.9/21.7/50.4/50.9 -
AusLaw-embedding RoC Aggregations Text Top-5 Citations 20.6/28.5/22.9/50.2/51.2 -
[…]
Results are formatted as setting1/2/3/4v1/4v2
Embeddings Index Granularity Query Output ACC@1 ACC@5
text-embedding-3-large RoC Aggregations Text Top-5 Citations 21.4/28.5/23.8/51.7/$52.1^{*}$
AusLaw-embedding RoC Aggregations Text Top-5 Citations 21.2/28.3/23.1/50.8/51.4 -
Table 1: Citation accuracy under various methods, including ours. For results marked by $\ddagger$, we used gpt-4o-mini due to cost. Best result is marked with . The rows are colour-coded for readability and easier comparison of settings referenced in the discussion §5.1. Text denotes ($S^{i-1}_{c_{k}},\texttt{Mask}(S^{i}_{c_{k}})$), see §4.1 for notations.